OpenStudy (calculusfunctions):

A train starts from rest and accelerates uniformly at 0.60 m/s² [E] until it reaches a velocity of 24 m/s [E]. It travels at 24 m/s [E] for a certain amount of time, then accelerates uniformly at -1.2 m/s² [E] until it comes to rest. The train's total displacement is 2400 m [E]. Find a). the distance that the train travelled at 24 m/s [E]. b). the total time taken for the entire trip.

OpenStudy (anonymous):

take it easy.., the answer is: a) the distance that the train travelled at 24 m/s is (s1) = 480 m b) the total time taken for the entire trip is 200 sekon, or 1/18 jam.., loot at the figure that i upload for you

OpenStudy (calculusfunctions):

Hi @gerryliyana thank you for trying!! I'm sorry to have to tell you that unfortunately your solution is incorrect. My students also had great difficulty with this problem although I think it's an excellent question to ask on a test. lol. I guess I wanted to prove to them that it is an easy question by posting it here and having other students solve it. I guess they were right considering that after eight days of posting, only one person out of many even attempted it. I commend you for that. I'm still hoping that others will try and that perhaps even you will redo the question again, with a different train of thought (pardon my pun). However if you choose to give up and wish to know the correct solution, let me know but I really hope you will persevere instead of quitting. I still haven't given up on my students either. Even though we have moved forward with a new unit, they're still encouraged to solve the problem. In fact some of them already have. Kudos to you and others who do not run from a challenge.

OpenStudy (nubeer):

ok first formula we need to remember.. a=v/t v=24, a =0.6 so t =40s this is the time it took him to reach the velocity of 24m/s lets suppose it moves with 24m/s for x secs.. and when the car start to decelrate at that point the time is "t" now its better to draw the velocity time graph which will make this question alot easier.  remember now t secs has been passed and its starting to declerate we can find the time it took to declerate.. v=24 , a =-1.2 as its decelrating we can take 1.2 so t = 20s ..

OpenStudy (nubeer):